# Cell divisions imprint long lasting elastic strain fields in epithelial tissues

### v1

@@ Acknowledgements
We would like to thank Michael Staddon, Frank Jülicher, Jonas Neipel for usefull discussions and comments. NAD, JFF, and GC acknowledge funding from the Max Planck Society, Excellence Cluster Physics of Life (EXC-2068-390729961), Deutsche Forschungsgemeinschaft (SPP1782/EA4/10-2), Deutsche Krebshilfe/MSNZ Dresden.
@@ V.1 Cell segmentation and tracking
The linear ablation experiments came from the pupal morphogenesis stage and were previously published in Ref. Piscitello-Gómez et al. (2023), where the authors analyzed initial retraction velocity of the ablated bonds. Here, we took these experiments and fully segmented and tracked cells in the entire field using the FIJI plugin, TissueAnalyzer Aigouy et al. (2016), using a bond cutoff of 2 pixels. Then, after the manual correction of segmentation errors, we use TissueMiner Etournay et al. (2016) to analyse changes in the cell shape and size.
The cell division experimental data came from the wing disc stage and was previously segmented, tracked and analyzed in Ref. Dye et al. (2017, 2021).
@@ V.2 Tissue relaxation after a laser ablation
After the ablation, nearby tissue relaxes to the new mechanical equilibrium. However, beyond $2$ minutes after the ablation we observed additional motion of the tissue which we interpret as the onset of the would healing response in the tissue. In this study, we are interested in the stress released by the ablation in the tissue, not the wound healing process, and therefore we limit our analysis to times up to $2$ minutes after the laser ablation.
@@ V.3 Tissue strain field measurement
Strain field consists of isotropic strain $U_{kk}(\mathbf{r},t)/2$ and shear strain $\tilde{U}_{ij}(\mathbf{r},t)$. To measure the strain field in the tissue we calculate the change in cell shape and size as described in Merkel et al. (2017); Duclut et al. (2021) at each timepoint. Since there are no cell rearrangements and divisions we estimate the accumulated tissue shear strain field in a cell from the difference between the cell shape tensors
@@
The only approximation is that we neglect to account for change in cell elongation due to local rotations. We determined the exact cell shear tensor at several timepoints and we found that the difference from our approximation are small. Therefore, we use the approximate expression for simplicity. We estimate the isotropic strain field in a cell from the change in cell area $a$
$U_{kk}(t)$ $=\log{\left(\frac{a(t)}{a(0)}\right)}\quad.$ (6)
We define the tissue strain field $U_{ij}(\mathbf{r},t)$ at position $\mathbf{r}$ to have the value of the strain tensor obtained for the cell inside of which point $\mathbf{r}$ lies.
@@ V.4 Discrete angular Fourier transform
To investigate the angular pattern of the strain field and study its radial profiles, we look at the angular discrete Fourier transformation of the tissue strain field. The strain in position $\bm{r}$ is given by ${U}_{ij}(r,\varphi)$ where $r$ is the distance to the center of the cut and $\varphi$ is the polar angle. We use the notation ${U}_{ij}(r,m)$ for the $m$-th angular mode of discrete angular Fourier transform at a distance $r$, and it is defined as,
@@
where $\varphi_{i}$s are $N$ uniformly distributed number between 00 and $2\pi$.
@@ V.5 Fit of the elastic theory to data
We measure force dipole, the ratio of shear modules to bulk modules, and viscosity by fitting the strain field pattern to the elastic and viscoelastic theories. First, we define a cost function
$E=\sum_{\alpha\in\rm{cells}}\Big{[}(U_{kk}^{\alpha}-u_{kk}(\bm{r}^{\alpha})^{2}+(\tilde{U}_{xx}^{\alpha}-\tilde{u}_{xx}(\bm{r}^{\alpha})^{2}+\\ (\tilde{U}_{xy}^{\alpha}-\tilde{u}_{xy}(\bm{r}^{\alpha})^{2}\Big{]},$ (8)
where ${U}_{ij}^{\alpha}$ is the cell strain measured from data, $\bm{r}^{\alpha}$ is the position of centre of each cell and ${u}_{ij}(\bm{r}^{\alpha})$ is the strain field predicted from theory in position $\bm{r}^{\alpha}$. By minimising $E$, we find the force magnitude of the force dipole, the ratio of the shear modulus to bulk modulus, and the relaxation timescales. In the elastic case, first, we minimise a steady state signal to find $\mu$ then we fix $\mu$ and repeat the minimisation process to find ${d}_{ij}$. In viscoelastic case, we use $\mu$ and ${d}_{ij}$ from the steady state elastic response, then we minimise $E$ at different time points to find $T_{ij}(t)$.
To analyse cell divisions, we first choose timepoint $t=25~{}\rm{min}$ and change the reference time $t_{0}$ from $-120$ to $-40$. We fit the data for different $t_{0}$ in that range and obtain corresponding set of $\mu(t_{0})$ values. We determine the parameter $\mu$ as the average $\langle\mu(t_{0})\rangle$ and we do not vary it in subsequent fitting of the strain field. Then, to determine the force dipole at each timepoint $t$ we fix $t_{0}=-60~{}\rm{min}$ fit the theory to the data to find the force dipole at different time points.
@@ V.6 Estimation of the tissue fluidization time
To measure the decay time scale of the anisotropic force dipole $\tilde{d}_{xx}$, we fit an exponential function to normalized force dipole $y=\tilde{d}_{xx}(t)/\langle a\rangle$ as $f(y)=\tilde{d}_{xx}(t=0)\exp[-t/\tau_{\alpha}]$ to observed data in Fig. 3 (D). From that, we obtain the fluidization times scale $\tau_{\alpha}=3.5\pm 0.7\rm{hr}$ and $\tilde{d}_{xx}(t=0)/\langle a\rangle=0.041\pm 0.001$.

### v2

@@ Acknowledgements
We would like to thank Michael Staddon, Frank Jülicher, Jonas Neipel for useful discussions and comments. Funding for the biological experiments was supported by the Max Planck Society and the Deutsche Forschungsgemeinschaft (Cluster of Excellence Physics of Life, EXC-2068-390729961 and SPP1782/EA4/10-2). NAD additionally acknowledges funding from Deutsche Krebshilfe/MSNZ Dresden. MP and ND acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project number DY 180/2-1 (eBer-24-23102) and PO 3023/2-1 - 544201605.
@@ V.1 Cell segmentation and tracking
The linear ablation experiments were performed during the pupal morphogenesis stage and were previously published in Ref. Piscitello-Gómez et al. (2023), where the authors analyzed the initial retraction velocity of the ablated bonds. Here, we re-analyzed movies of those experiments, and fully segmented and tracked cells in the entire field using the FIJI plugin, TissueAnalyzer Aigouy et al. (2016), using a bond cutoff of 2 pixels. Then, after the manual correction of segmentation errors, we use TissueMiner Etournay et al. (2016) to analyze changes in the cell shape and size.
The cell division experimental data came from the wing disc stage and was previously segmented, tracked, and analyzed in Ref. Dye et al. (2017, 2021).
@@ V.2 Tissue relaxation after a laser ablation
After the ablation, nearby tissue relaxes to a new mechanical equilibrium. In Experiment 1 reported in the main text, we estimate equilibrium to be reached after about $30s$ after ablation, when the tissue does not visibly evolve any more. However, this equilibrium state is not permanent, as at about $1$ minute after the ablation, we observe additional motion of the tissue, which we interpret as the onset of the wound-healing response. In this work, we are interested in the stress response to the ablation, not the wound-healing process, and therefore we limit our analysis to times below $1$ minute after the laser ablation. Note that in Experiments 2-4, reported in detail in SI, the additional patterns appear to emerge earlier and we could not extract as clear an intermediate regime as in Experiment 1. However, even in these experiments the profiles of strain observed during the relaxation are consistent with the linear elastic theory, thereby further strengthening our results from Experiment 1.
@@ V.3 Tissue strain field measurement
Strain field consists of isotropic strain $U_{kk}(\mathbf{r},t)/2$ and shear strain $\tilde{U}_{ij}(\mathbf{r},t)$. To determine the strain field from the changes of cell shape and area, we use the method based on the cell triangulation Etournay et al. (2015); Merkel et al. (2017); Duclut et al. (2021) in which the co-rotational change of the cell elongation tensor $Q_{ij}$ corresponds to the tissue shear rate, and changes in the logarithm of cell area to the isotropic strain rate. Here, we approximate the accumulated tissue strain by the change in cell elongation tensor and in the logarithm of cell area for pure shear and isotropic components, respectively. To define the cell elongation tensor $Q_{ij}$, we choose the sub-cellular triangulation of the tissue obtained by connecting two adjacent vertices to the geometric center of cells, as described in Duclut et al. (2021). We estimate the accumulated tissue shear strain field in a cell from the difference between the cell shape tensors
@@
The only approximation is that we neglect to account for changes in cell elongation due to local rotations. We determined the exact cell shear tensor at several timepoints, and we found that the differences from our approximation are small. Therefore, we use the approximate expression for simplicity. The isotropic strain field in a cell is defined as the accumulated relative area change rate Etournay et al. (2015); Merkel et al. (2017), which we measure relative to the initial cell state as
$U_{kk}(t)$ $=\log{\left(\frac{a(t)}{a(0)}\right)}\quad,$ (6)
where $a$ is the cell area.
Finally, we define the tissue strain field $U_{ij}(\mathbf{r},t)$ at position $\mathbf{r}$ to have the value of the strain tensor obtained for the cell inside of which point $\mathbf{r}$ lies.
@@ V.4 Discrete angular Fourier transform
We examine the angular discrete Fourier transformation of the tissue strain field to investigate its angular pattern and study its radial profiles. The strain in position $\bm{r}$ is represented by ${U}_{ij}(r,\varphi)$, where $r$ is the distance to the center of the cut and $\varphi$ is the polar angle. We use the notation ${U}_{ij}(r,m)$ for the $m$-th angular mode of discrete angular Fourier transform at a distance $r$, and it is defined as,
@@
where $\varphi_{i}$ are $N$ uniformly distributed numbers between 00 and $2\pi$. In rare cases a point with coordinates $(r,\varphi_{i})$ does not belong to a cell, making ${U}_{ij}(r,\varphi)$ undefined at that point. To handle such cases, we modify the Fourier transform as follows,
${U}_{ij}(r,m)=\frac{\sum_{i=1}^{N}{U}_{ij}(r,\varphi_{i})e^{im\varphi_{i}}\iota(r,\varphi_{i})}{\sum_{i=1}^{N}\iota(r,\varphi_{i})},$ (8)
where $\iota(r,\varphi_{i})=1$ if the point belongs to a cell and 00 otherwise.
[…]
@@ V.5 Fit of the elastic theory to data
We measure the force dipole, the ratio of shear moduli to bulk moduli, and viscosity by fitting the strain field pattern to the elastic and viscoelastic theories. First, we define a cost function
$E=\sum_{\alpha\in\rm{cells}}\Big{[}(U_{kk}^{\alpha}-u_{kk}(\bm{r}^{\alpha}))^{2}+(\tilde{U}_{xx}^{\alpha}-\tilde{u}_{xx}(\bm{r}^{\alpha}))^{2}+\\ (\tilde{U}_{xy}^{\alpha}-\tilde{u}_{xy}(\bm{r}^{\alpha}))^{2}\Big{]},$ (9)
where ${U}_{ij}^{\alpha}$ is the cell strain measured from data, $\bm{r}^{\alpha}$ is the position of the center of each cell and ${u}_{ij}(\bm{r}^{\alpha})$ is the strain field predicted from theory at position $\bm{r}^{\alpha}$. By minimizing $E$, we find the magnitude of the force dipole, the ratio of the shear modulus to bulk modulus, and when fitting the relaxation dynamics, the relaxation timescales.
When fitting the steady state strain field measured in Experiment 1, we wanted to take into account the fluctuations in the steady state. For this, we sampled the strain field at 5 timepoints uniformly spaced between $36s$ and $58.5s$ after ablation. We constructed datasets by bootstrapping the sampled data $10$ times, and we fit the linear elastic theory for each dataset. The reported parameter values and uncertainties are mean values and standard deviations of these 10 fit outcomes.
To fit the relaxation dynamics and extract the relaxation timescale, we consider the strain field at each time point $t_{i}$, spaced $4.5s$ apart, during the relaxation and use the steady state parameters and force dipole to fit the elements of the matrix $\bm{T}(t_{i})$. Finally, we fit an exponential function to the obtained matrix element values at different time points, Fig. 2.C.ii, and report the relaxation timescale values and their uncertainties obtained by this exponential fit.
To analyze cell divisions, we first reduce noise in the strain field by applying a moving average over a $10$-minute window. Next, we select the time point $t=30\rm{min}$ and systematically vary the reference time $t_{0}$ from $-120\rm{min}$ to $-40\rm{min}$. For each $t_{0}$ within this range, we fit the steady state linear elastic theory to the measured strain field to extract the corresponding set of $\mu(t_{0})$ values. The parameter $\mu$ is then determined as the average $\langle\mu(t_{0})\rangle$, which remains fixed for the subsequent strain field analyses (see SI Fig. 14). This is done to more precisely measure the parameter $\mu$, which cannot be extracted independently of $d_{0}$ when traceless symmetric components of the force dipole vanish. Namely, in that case it is impossible to factorize the product $\mu D_{0}$ as the two factors only appear in the form of the product, see SI Eqs. 9 and 10. Finally, to compute the force dipole at each timepoint $t$ we set $t_{0}=-60\rm{min}$, and we perform the fit at each time point using the bootstrapping algorithm over the ensemble of all divisions, obtaining the force dipole values and their uncertainties at each time point, while using the previously obtained values of $\mu$ as a constant.
@@ V.6 Estimation of the tissue fluidization time
To measure the decay timescale of the anisotropic force dipole $\tilde{d}_{xx}$, we fit an exponential function to normalized force dipole $y=\tilde{d}_{xx}(t)/\langle a\rangle$ as $f(y)=\tilde{d}_{xx}(t=0)\exp[-t/\tau_{\alpha}]$ to observed data in SI Fig. 16. From that, we obtain the fluidization timescale $\tau_{\alpha}=3.5\pm 0.7\rm{hr}$ and $\tilde{d}_{xx}(t=0)/\langle a\rangle=0.53\pm 0.01$. The reported errors correspond to the range of parameters over which an exponential function successfully fits the data.